OpenStudy (anonymous):

the ratio of two numbers is 5/4 and the first number is 7 more than the second number. to determine each of these numbers, let x be the first number. Since the first number is 7 more than the second number, and since the first number is x, what is the second number?

OpenStudy (anonymous):

Let first number be X. Second number = X - 7 (first is 7 bigger than second) Ratio of two numbers = Bigger number / Smaller Number BIgger number is the first number. Ratio of two numbers = (x) / (x-7) = 5/4 By simplifying: 4x = 5(x-7) x=? Thus, x-7 = ?

OpenStudy (anonymous):

wouldn't you cross multiply the 5(x-7) ?

OpenStudy (anonymous):

idk if that's what its called

OpenStudy (anonymous):

4x = 5(x-7) 4x = 5x-35 35 = x x-7 = ?

OpenStudy (anonymous):

28